OpenStudy (anonymous):

CHECK MY ANSWER PLZ. I SAY IT"S C :) Which of the following is a solution for the system y = -3x + 5 and y = 2x + 5? A. 1,7 5,0 2,-1 0,5

OpenStudy (anonymous):

Did you substitute to check your answer?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

the awser is we

OpenStudy (anonymous):

if u worked it out and got C then it is C

OpenStudy (anonymous):

I just need some one to verify.... :/

OpenStudy (anonymous):

\[-1=-3\times2+5\] Is that correct @GodGirl?

OpenStudy (anonymous):

yes. @AMYCARTER

OpenStudy (anonymous):

But, its not the correct solution for the other equation.

OpenStudy (anonymous):

Yep you're right @GodGirl so it can't be C.

OpenStudy (anonymous):

try A

OpenStudy (anonymous):

@AMYCARTER Thanks.

OpenStudy (anonymous):

@kylefoster58 I tried all of them.

OpenStudy (anonymous):

No problem, do you still need more help? Did you find the answer?

OpenStudy (anonymous):

so wich one is it?

OpenStudy (anonymous):

witch

OpenStudy (anonymous):

@AMYCARTER I already had the answer. I just needed someone to check it. @kylefoster58 Its C (-1,2)

OpenStudy (anonymous):

but it doesnt go with the second one i thought

OpenStudy (anonymous):

I thought you said it doesn't work for the second equation?

OpenStudy (anonymous):

It doesn't.

OpenStudy (anonymous):

But, there cant be 2 answers...

OpenStudy (anonymous):

then that isn't it!

OpenStudy (anonymous):

the answer is not C

OpenStudy (anonymous):

anyone wanna give me a medal? i so close to getting human calculator!

OpenStudy (anonymous):

A isn't right either @GodGirl so it could be B or C, want to try?

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

B or D

OpenStudy (anonymous):

Oh yeah B or D that was a typo

OpenStudy (anonymous):

ok, im working it out on paper @AMYCARTER

OpenStudy (anonymous):

yea.

OpenStudy (anonymous):

\[5=-3\times0+5\] Is this right @GodGirl?

OpenStudy (anonymous):

@AMYCARTER wanna give me a medal? i am almost to human calculator

OpenStudy (anonymous):

A works for the 2nd equation @AMYCARTER

OpenStudy (anonymous):

yes @AMYCARTER

OpenStudy (anonymous):

Yes I know but you need to find one solution that works for BOTH equations okay?

OpenStudy (anonymous):

Yeah ik I was just telling you... .-.

OpenStudy (anonymous):

Now let's try A for the second equation, is this right @GodGirl \[5=2\times0+5\]

OpenStudy (anonymous):

yas

OpenStudy (anonymous):

yess @AMYCARTER

OpenStudy (anonymous):

Thats not A. Thats B

OpenStudy (anonymous):

A is (1,7)

OpenStudy (anonymous):

Well looks like B is your answer @GodGirl.

OpenStudy (anonymous):

This is what I got when trying B for equation 1. (5,0) (x,y) y=-3x+5 0=-3(5)+5 0=-10

OpenStudy (anonymous):

@AMYCARTER

OpenStudy (anonymous):

You were substituting it wrong in the equation @GodGirl.

OpenStudy (anonymous):

Oh I see. But, either way the 0 is first in the equation. i think the answer is actually D

OpenStudy (anonymous):

Did you try it, maybe I didn't see it right or someting

OpenStudy (anonymous):

*something

OpenStudy (anonymous):

Yeah I got it. y=-3x+5 y=2x+5 5=-3(0)+5 5= 2(0)+5 5=5 5=5 @AMYCARTER

OpenStudy (anonymous):

(0,5)